OpenStudy (anonymous):

what are the solutions of x^2+4x+5=0

OpenStudy (anonymous):

x^2+4 x+5 = 0 Subtract 5 from both sides: x^2+4 x = -5 Add 4 to both sides: x^2+4 x+4 = -1 Factor the left hand side: (x+2)^2 = -1 Take the square root of both sides: abs(x+2) = i Eliminate the absolute value: x+2 = -i or x+2 = i Subtract 2 from both sides: x = -2-i or x+2 = i Subtract 2 from both sides: x = -2-i or x = -2+i

OpenStudy (anonymous):

\[x=\frac{-b+/-\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[x=\frac{-4+/-\sqrt{4^2-4(1)(5)}}{2}\]

OpenStudy (anonymous):

\[x=\frac{-4+/-4i}{2}=-2+/-2i\]